OpenStudy (anonymous):

Please HELP: Find the equation of the parabola. GIVEN: focus (-6, 0) ; Directrix x=6

OpenStudy (zehanz):

If the focus is \[F(\frac{ 1 }{ 2 }p,0)\]and the directrix is\[x=-\frac{ 1 }{ 2 }p\]then the vertex form of the equation is\[y^2=2px\]The given focus and directrix gives: p = -12. The equation must be:\[y^2=-24x\]